Determine the equation of the line that is perpendicular to the y=-1/5 x and passes through the point (1,9)
Answer by stanbon(75887)

Determine the equation of the line that is perpendicular to the y=-1/5 x and passes through the point (1,9)
--------------
The given equation has a slope of (1/5)
Any perpendicular line must have slope = -5
---------
The equation form will be y = mx+b
Substitute to find "b":
9 = -5*1+b
b = 14
--------
EQUATION:
y = -5x+14
================
Cheers,
Stan H.